OpenStudy (tparker22):

Which of the following points lies on the line whose equation is y = 2x - 3? I think it's (-1, 4) can you please verify this for me so I can know if I'm right? Medal being given!!

jimthompson5910 (jim_thompson5910):

Let's find out. (-1,4) means x = -1 and y = 4. So let's replace x with -1 and y with 4 ------------------ y = 2x - 3 y = 2(-1) - 3 ... replace x with -1 4 = 2(-1) - 3 ... replace y with 4; now simplify 4 = -2 - 3 4 = -5 ... false the last equation is false, so y = 2x - 3 is false when (x,y) = (-1,4) This means (-1,4) is NOT on the line y = 2x - 3

OpenStudy (tparker22):

Okay, so would (0, 3, be it?

jimthompson5910 (jim_thompson5910):

plug in (x,y) = (0,3) and see if you get a true equation or not

jimthompson5910 (jim_thompson5910):

y = 2x - 3 3 = 2(0) - 3 .. plug in (x,y) = (0,3) 3 = 0 - 3 3 = -3 ... false So (0,3) is NOT on the line

OpenStudy (tparker22):

So that means I'm left with (1, -1) :)

jimthompson5910 (jim_thompson5910):

try it out